omework Assignment Saved Help Saw Exercise 12-2 (LO12-1) Wha t is the critical Fvalue when the sample size for the numerator is four and the sample size for the denominator is seven? Use a e-tailed test and the 0.01 significance level. (Round your answer to 2 decimal places.)
Answer #1

Answer

We have sample size n1 = 4 for numerator and sample size n2 = 7 for denominator

first, we will find the degree of freedom. We know that degree of freedom = n-1

so, degree of freedom for numerator = 4-1 = 3 and degree of freedom for denominator = 7-1 =6

Now, we have to find the 3 in first row in F critical values table and 6 in the first column, then select the intersection of both(keep 0.01 level when finding intersection).

we get, F critical = 9.78

So, required F critical value is 9.78 (rounded to 2 decimals)